\begin{thebibliography}{xx}
\bibitem{BB} Bertrand, J.; Bertrand, P.
Symbolic calculus on the time-frequency half-plane. 
J. Math. Phys. 39, No.8, 4071-4090 (1998). physics/9708027
\bibitem{B} Bieliavsky, P. Strict Quantization of Solvable Symmetric Spaces.
math.QA/0010004, to appear in the Journal of Symplectic Geometry.
\bibitem{BM} Bieliavsky, P.; Massar, M. Oscillatory integral 
formulae for left invariant star product on a class of Lie groups. Letter in Math. Phys.
\bibitem{G} Guelfand, I.M.; Chilov, G.E.
Les distributions. Tome II: Espaces fondamentaux. Traduit du russe par S.
Vasilach (French)
Collection universitaire de mathematiques. 15. Paris: Dunod. (1964). 
\bibitem{GZ} Giaquinto, A.; Zhang, J.J.
Bialgebra actions, twists, and universal deformation formulas. (English)
J. Pure Appl. Algebra 128, No.2, 133-151 (1998).
\bibitem{H} Huynh, T. V. Invariant $\star$-quantization associated with 
the affine group. J. Math. Phys. {\bf 23}(6), 1082-1087 (1982).
\bibitem{M} Maillard, J-M.
Sur le produit de convolution gauche et la transformée de Weyl des
distributions tempérées. (On the twisted convolution product and the Weyl
transform of the tempered distributions). (French)
C. R. Acad. Sci., Paris, Sér. I 298, 35-38 (1984).
\bibitem{OMMY} Omori, H.; Maeda, Y.; Miyazaki, N.; Yoshioka,
A.
Anomalous quadratic exponentials in the star-products. (English)
RIMS Kokyuroku 1150, 128-132 (2000).
\bibitem{R} Rieffel, M.A.
Deformation quantization for actions of $\R^d$. (English)
Mem. Am. Math. Soc. 506, 93 p. (1993).
\bibitem{S} Schwartz, L.
Théorie des distributions. (Distribution theory). Nouveau tirage. (French)
Paris: Hermann. xii, 420 p. (1998).
\end{thebibliography}

\documentclass[numreferences]{kluwer}
\usepackage{amssymb}
\usepackage{latexsym}

\newtheorem{thm}{Theorem}[section] 
\newtheorem{dfn}[thm]{Definition}
\newtheorem{rmk}[thm]{Remark}

\newtheorem{prop}[thm]{Proposition}
\newtheorem{lem}[thm]{Lemma}

\newcommand{\Pf}{{\em Proof}. }
\newcommand{\EPf}
{%
\mbox{}%
\nolinebreak%
\hfill%
\rule{2mm}{2mm}%
\medbreak%
\par%
}

\newcommand{\C}{\mathbb C} 
\newcommand{\R}{\mathbb R}

\newcommand{\g}{{\cal G}{}}

\newcommand{\s}{{\cal S}{}} 

\renewcommand{\d}{{\cal D}{}}

\renewcommand{\o}{{\cal O}{}}

\renewcommand{\Re}{\mbox{Re}}
\newcommand{\sm}{\star_{\nu}^{M}}
\newcommand{\sw}{\star_{q}^{W}}
\newcommand{\sn}{\star_{\nu}}
\newcommand{\skq}{\star_{q}^{(k)}}

\newcommand{\ronu}{\rho_{\nu}}
\newcommand{\ronuhat}{{\hat{\rho}}_{\nu}}
\newcommand{\LT}{{\cal L}{}}
\newcommand{\FT}{{\cal F}{}}
\newcommand{\ZT}{{\cal Z}_\nu{}}
\newcommand{\taukq}{\tau^{(k)}_q{}}
\newcommand{\Tkq}{T^{(k)}_q{}}

\def\Der{{\cal	D}er}

\newcommand{\sab}{{\cal S}_\alpha^\beta{}}
\newcommand{\sba}{{\cal S}_\beta^\alpha{}}
\newcommand{\sA}{{\cal S}^{\sigma(\alpha_{1},\alpha_{2})}_{(\alpha_{1},\alpha_{2})}(2){}}
\newcommand{\Ekab}{E^{(k)}_{(\alpha_{1},\alpha_{2})}{}}

\title{Convergent star product algebras on ``$ax+b$''}
\author{{\bf Pierre Bieliavsky}\\
Universit\'e Libre de Bruxelles, Belgium\\
e-mail: pbiel@ulb.ac.be\\
{\bf Yoshiaki Maeda}\\
Keio University, Japan\\
e-mail: maeda@math.keio.ac.jp}
\begin{document}
\maketitle
\section{Introduction}
The notion of convergent star product is generally understood as the 
data of a one parameter family $\{E_t\}_{t\in I}\subset C^\infty(M)$ of 
function algebras on a Poisson manifold $(M,\{\, , \,\})$.  On each of 
them
one is given an associative algebra 
structure $\star_t$ which respect to which the function space $E_t$ is 
closed. The family of products $\{\star_t\}$ should moreover define in some 
sense a deformation of the commutative pointwise product of functions in 
the direction of the Poisson structure $\{\, , \,\}$.

Stable function algebras for the Weyl-Moyal product have been studied in 
various contexts. For instance, see:
\begin{enumerate}
\item[-] \cite{M} for such a study in the framework of tempered 
distributions on a symplectic vector space;
\item[-] \cite{R} for a $C^\star$-algebraic study on 
$\R^d$-manifolds;
\item[-] \cite{OMMY} for non-tempered stable function spaces on $\C^n$.
\end{enumerate}
A special feature of the Weyl-Moyal star product---independently of 
the functional framework--- is its maximal invariance under the group 
of affine transformations with respect to a {\sl flat} affine 
connection. This can be rephrased by saying that the Weyl-Moyal 
quantization is universal with respect to actions of $\R^d$  
\cite{R}. A natural question is then the one of defining universal 
(convergent) deformations for non-Abelian Lie group actions. 
In the formal framework, this has been investigated in \cite{GZ}.
In, \cite{BB,H}, such formulae in the case of $ax+b$ have been studied within the context of
Wigner formalism and signal analysis. However, the question of defining
an adapted functional framework has not been investigated. In 
\cite{B,BM}, one finds a functional analytic study for solvable Lie 
group actions and symmetric spaces in the tempered distributions and/or $C^\star$-context.

It therefore appears quite challenging to investigate the problem of 
defining non-tempered (e.g. exponential growth) function spaces on 
such a non-Abelian Lie group which are stable under some left-invariant 
(convergent) star product. In other words, studying a situation where 
non-temperedness and non-linear invariance mix. This is what is done 
in this paper for the particular case of the group $ax+b$. More 
precisely, we first start by giving a construction of a left-invariant 
star product on the (symplectic) group manifold underlying the Lie 
group $ax+b$.  This star product is obtained via an equivalence 
transformation $T$ performed on Moyal's product (which is not 
left-invariant). The equivalence $T$ involves two ingredients: a 
partial Laplace transform and a family 
$\{\phi_{\nu,\gamma}:\C\to\C\}_{\nu,\gamma\in\C}$ of holomorphic maps.
For special values of of the parameters $\nu$ and $\gamma$ one refinds 
the functional calculus studied in \cite{B,BM}. But for other values, 
one can define stable function algebras constituted by type-S functions.
Using a holomorphic presentation of 
these spaces, one gets non-commutative algebra structures on spaces of
entire functions on $\C^2$. Such a space contains functions of 
exponential growth, one therefore has a non-tempered invariant 
calculus.

\noindent{\bf Acknowledgment} We thank Daniel Sternheimer for having 
communicated important references.
\section{Formal star products on ``$ax+b$''}\label{FORMAL}
In this section, we briefly recall results appearing in \cite{B,BM}.
Let $\g$ denote the Lie algebra of the group of affine 
transformations of the real line. Formally, one has 
$\g=\mbox{span}_{\R}\{A,E\}$ with table $[A,E]=2E$. Consider 
the linear map $\lambda:\g\to C^\infty(\R^2):X\mapsto\lambda_X$
defined by $\lambda_A(a,l)=2l;\quad\lambda_E(a,l)=e^{-2a}$, where 
$\R^2=\{(a,l)\}$. One then checks that the map $\lambda$ is a homomorphism 
of Lie algebras when $C^\infty(\R^2)$ is endowed with the symplectic Poisson 
bracket $\{\, , \,\}:=\partial_a\wedge\partial_l$. Moreover, if $\sm$ 
denotes the formal Moyal star product on $C^\infty(\R^2)[[\nu]]$ 
(i.e 
$u\sm v=u\,\exp{(\nu\stackrel{\leftarrow}{\partial_a}\wedge
\stackrel{\rightarrow}{\partial_l})}v\qquad u,v\in C^\infty(\R^2)[[\nu]]$),
one has $\left[\lambda_A,\lambda_E\right]_\nu=2\nu\{\lambda_A,\lambda_E\}$
(where $\left[u,v\right]_\nu:=u\sm v-v\sm u$). In particular, the formula
$$
\ronu(X)u:=\frac{1}{2\nu}\left[\lambda_X,u\right]_\nu\qquad X\in\g,u\in C^\infty(\R^2)[[\nu]]
$$
defines a homomorphism of Lie algebras 
$$
\ronu:\g\to\Der(C^\infty(\R^2)[[\nu]],\sm).
$$
Explicitly, one has 
$\ronu(A)u=-\partial_au;\quad\ronu(E)u=-\frac{e^{-2a}}{\nu}\sinh(\nu\partial_l)u$.
Intertwining the representation $\ronu$ by a transformation of the type
$$
\LT(u)(a,z):=\int_\R e^{-zl}u(a,l)\,dl,
$$
one gets
$$
\begin{array}{ccc}
\ronuhat(A)\LT(u):=\LT(\ronu(A)u) & = & -\partial_a\LT(u);\\
\ronuhat(E)\LT(u):=\LT(\ronu(E)u) & = & -\frac{e^{-2a}}{\nu}\sinh(\nu 
z)\LT(u),
\end{array}
$$
where we assumed $u(a,\pm\infty)=0$. Now, set formally
$$
\ZT(u)(a,z):=\int_\R \exp{\left(\gamma\frac{1}{\nu}\sinh(\nu z)l\right)}\,u(a,l)\,dl,
$$
and 
$$
f\bullet_\nu g:=\ZT(\ZT^{-1}f.\ZT^{-1}g)\qquad(\gamma\in\C_{0}).
$$
\begin{prop}
For all $X\in\g$, $\ronuhat(X)$ is a derivation of the commutative 
product $\bullet_\nu$.
\end{prop}
In other words, the associative formal product $u\sn v:=T^{-1}(Tu\sm Tv)$ 
where $T=\LT^{-1}\ZT$ is invariant under the infinitesimal action 
$\g\to\Gamma T(\R^2):X\mapsto{}^\sharp(d\lambda_X)$ where ${}^\sharp(d\lambda_X)$ 
denotes the Hamiltonian vector field associated (via the symplectic 
structure) to the function $\lambda_X$. This action of $\g$ turns out to 
exponentiate as a global simply transitive symplectic action of the group $G=``ax+b"$ on 
$\R^2$, providing an identification of the group manifold 
underlying $G$ with $\R^{2}$.

The integral form of the transformation $\ZT$ allows to define specific 
functions algebras on $\R^2$ (as opposed to power series  algebras) stable 
under the product $\sn$ where the formal Moyal product $\sm$ is replaced 
by its ``convergent" version: the Weyl product. The case where $\nu\in 
i\R,\quad\gamma=1,\quad z\in i\R$ has been studied in \cite{BM,B}. In 
what follows, we are concerned with the general case where $\nu\in 
i\R,\quad\gamma\in U(1),\quad z\in \C$. We end up this section by observing 
that the intertwiner $T=\LT^{-1}\ZT$ can be expressed as 
$$
T=\LT^{-1}\circ(\phi_{\nu,\gamma})^*\circ\LT
$$
where we set $\phi_{\nu,\gamma}(z):=\frac{\gamma}{\nu}\sinh(\nu z)$. The 
map $\phi_{\nu,\gamma}$ will be referred in the sequel as the ``twisting 
map" (see Section \ref{TM}).
\begin{rmk}
{\rm
An alternative simple way for obtaining an explicit formula of an 
invariant star product on $ax+b$ is based on the following observation
\cite{BB,H}. The symplectic group manifold underlying $ax+b$ can be 
seen as an open coadjoint orbit $\cal O$ in $\g^{\star}$. Quantizing 
the 
Poisson manifold $\g^{\star}$ via the universal enveloping algebra 
product and  then restricting to $\cal O$ yields an invariant star 
product on $\cal O$ hence a left-invariant one on $ax+b$. It is classical 
that an oscillatory integral formula for this product can be written 
down in terms of the Campbell-Baker-Haussdorff function (see 
\cite{BB} for explicit computation: Formulae 5.8 and 5.12). Our product
$\star_{\nu}$ described above is different from the universal 
enveloping algebra product. Indeed, their invariance diffeomorphism 
groups do not coincide \cite{B}.
}
\end{rmk}
\section{Fundamental spaces of type ${\cal S}$}\label{TYPE-S}
In this section, we follow Chapter IV of I.M.Guelfand's book \cite{G}.
We will denote by $\o(\C^m)$ the space of holomorphic (entire) functions 
on $\C^m$. 
\begin{dfn} Let $\alpha,\beta\in\R^m$.
The fundamental space $\sab(m)$ is defined as the space of holomorphic 
functions $\varphi\in\o(\C^m)$ such that there exists $a,b\in(\R^+)^m$ 
and $C>0$ with
$$
|\varphi(x+iy)|\leq 
C\,\exp\left(-a|x|^{\frac{1}{\alpha}}+b|y|^{\frac{1}{1-\beta}}\right),
$$
where we adopt the usual notations~:
$a|x|^e=\sum_ja_j|x_j|^{e_j}\quad(a,x,e\in\R^m); 
\,\frac{1}{\alpha}=(\frac{1}{\alpha_1},...,\frac{1}{\alpha_m});\, 
1-\beta=(1-\beta_1,...,1-\beta_m)$.
\end{dfn}
Every element $\varphi\in \sab(m)$ is entirely determined by its restriction 
the ``real axis" $\varphi(x)\quad x\in\R^m$. We will often identify 
the space $\sab(m)$ with the subspace $\left(\sab(m)\right)_x$ of 
$C^\infty(\R^m)$ constituted by the restrictions. In order to consider 
only non-trivial spaces, we will assume 
$\alpha+\beta\geq1;\quad\alpha>0;\quad\beta>0$. We will denote by 
$\FT(u)(\xi)$ the Fourier transform of the function $u\in L^1(\R^m)$~:
$$
\FT(u)(\xi):=\int_{\R^m}e^{i\xi.x}u(x)\, dx ,
$$
where $\xi.x$ denotes the canonical dot product on $\R^m$. For even $m=2n$, 
we will denote by $J$ the endomorphism of $\R^{2n}$ defined by the matrix
$$
[J]:=\left(
\begin{array}{cc}
0 & I_n\\
-I_n & 0
\end{array}\right)
$$
where $I_n$ is the $n\times n$ identity matrix. We denote by $\omega$ the 
bilinear symplectic structure on $\R^{2n}$ defined by $\omega(x,y):=x.Jy$.
\begin{dfn}
We define the symplectic Fourier transform of the function $u\in 
L^1(\R^{2n})$  as 
$$
S\FT(u)(y):=\int_{\R^{2n}}e^{i\omega(x,y)} u(x)\, dx \qquad (y\in\R^{2n}).
$$
\end{dfn}
Equivalently, one has $S\FT=J^*\circ\FT$, which yields (see \cite{G})
\begin{lem}
One has $$S\FT(\sab(2n))=\s^{\sigma(\alpha)}_{\sigma(\beta)}(2n),$$ where for 
$\alpha=(\alpha_1,\alpha_{2})\in\R^{{2n}}\,(\alpha_{i}\in\R^n)$, we set
$\sigma(\alpha):=(\alpha_2,\alpha_1)$.
\end{lem}
\begin{dfn}
For $u,v\in L^1(\R^{2n})$, one defines their twisted convolution by 
$$
u\times_qv(x):=\int_{\R^{2n}}e^{iq\omega(x,y)}u(y)\, 
v(x-y)\,dy\qquad(q\in\R_0).
$$
\end{dfn}
\begin{lem}\label{MULT}
One has $$\sab(m).\s^{\beta'}_{\alpha'}(m)\subset
\s^{\mbox{\rm max}(\beta,\beta')}_{\mbox{\rm min}(\alpha,\alpha')}(m);$$ where we set
$\mbox{\rm max}(\beta,\beta'):=(\mbox{\rm 
max}(\beta_1,\beta'_1),...,\mbox{\rm max}(\beta_m,\beta'_m))$.
\end{lem}
\Pf
For simplicity, we assume $m=1$. Let $\varphi\in\sab$ and $\varphi'\in\s^{\beta'}_{\alpha'}$.
Then, for $z=x+iy$, one has
{\small{
\begin{eqnarray*}
|\varphi\varphi'(z)|\leq CC'\exp\left(
-a|x|^{\frac{1}{\alpha}}+b|y|^{\frac{1}{1-\beta}}-a'|x|^{\frac{1}{\alpha'}}+
b'|y|^{\frac{1}{1-\beta'}}\right)
\end{eqnarray*}
}}
which is lower than 
{\small{
$C"\exp\left(
-a"|x|^{\mbox{\rm max}(\frac{1}{\alpha},\frac{1}{\alpha'})}
+b"|y|^{\mbox{\rm 
max}(\frac{1}{1-\beta},\frac{1}{1-\beta'})}\right)$}} for 
some $C",a",b"$. But, 
$\mbox{max}(\frac{1}{\alpha},\frac{1}{\alpha'})=\frac{1}{\mbox{\rm min}({\alpha},{\alpha'})}$ and 
$\mbox{max}(\frac{1}{1-\beta},\frac{1}{1-\beta'})=\frac{1}{1-\mbox{max}(\beta,\beta')}$.
\EPf
\begin{lem}
Let $u,v\in\sab(2n)$. Then
$u\times_qv\in\s^{\sigma{\alpha}}_{\sigma{\beta}}(2n)$.
\end{lem}
\Pf
Changing the variables following $y\mapsto-y$, one gets 
$u\times_qv=\left[d_q^* S\FT(\tilde{u}\alpha_xv)\right]^{\tilde{}}$ where 
$\tilde{u}(x):=u(-x)$, $(\alpha_xv)(y):=v(y+x)$ and where $d_q$ denotes the 
dilation in $\R^m$~: $x\mapsto qx \quad(q\in\R)$. Hence $u\times_qv\in 
S\FT\left(\s^{\mbox{\rm max}(\beta,\beta)}_{\mbox{\rm 
min}({\alpha},{\alpha})}(2n)\right)=\s^{\sigma(\alpha)}_{\sigma(\beta)}(2n)$.
\EPf
\begin{dfn}(see e.g. \cite{M})
The Weyl product between $u$ and $v$ in $L^{1}(\R^{2n})$ is defined by
$$
u\sw v:=S\FT\left[S\FT(u)\times_qS\FT(v)\right].
$$
\end{dfn}
\begin{prop}\label{WEYL}\cite{OMMY}
Let $u,v\in\sab(2n)$, Then $u\sw v\in\s_{\sigma(\beta)}^{\sigma(\alpha)}(2n)$. 
In particular, the space $\s_\alpha^{\sigma(\alpha)}(2n)$ is stable under 
the Weyl product.
\end{prop}
\begin{rmk}
{\rm
The space $\s_\alpha^{\sigma(\alpha)}(2n)$ is stable under the pointwise 
multiplication as well.
}
\end{rmk}
\section{Laplace Transformation}
In this section we follow L. Schwartz' book \cite{S}. We adopt the 
following notations. We denote by $\d$ the space of compactly supported 
smooth functions on $\R$ endowed with the topology of test 
functions. We denote by $\d'$ the space of distributions on $\R$.
Also, if $\Omega$ is an open domain in $\C$, we set $\o(\Omega)$ for the 
space of holomorphic functions on $\Omega$.
\begin{dfn}Let $\Gamma$ be an open interval in $\R$.
The fundamental space $\s'_\Gamma$ is defined as the space of 
distributions $T\in\d'$ such that for all $\xi\in\Gamma$, the distribution 
$\exp(-\xi.x)T_x$ is tempered. We denote by $\o_\Gamma$ the space of 
holomorphic functions $F\in\o(\Gamma+i\R)$ such that for all compact set 
$K\subset\Gamma$, the restriction $F|_{K+i\R}(\xi+i\eta)$ is bounded by a 
polynomial in $\eta$.
\end{dfn}
\begin{prop}
One defines the Laplace transform of an element $T\in\s'_\Gamma$ as the Fourier 
transform of $e^{-\xi.x}T_x$
$$
\LT(T)(\xi+i\eta):=\left(\FT_x\left[\exp(-\xi.x)T_x\right]\right)(\eta).
$$
Then, setting $z=\xi+i\eta\in\Gamma+i\R$, one has a linear isomorphism
$$
\LT:\s'_\Gamma\to\o_\Gamma.
$$
\end{prop}
\begin{rmk}
{\rm
Provided the following integrals make sense, one has
$$
\LT(T)(z)=\int_\R e^{-zx}T_x\,dx\mbox{ and }
\LT^{-1}F(x)=\int_{c+i\R}e^{zx}F(z)dz,
$$
where $c$ is any element of $\Gamma$. Indeed, if $F=\LT T$ one has for all 
$c\in\Gamma$~:
{\small{
\begin{eqnarray*}
\int_{c+i\R}e^{zx}F(z)dz=e^{xc}\int_\R F(c+it) e^{ixt}dt=\\
=
e^{xc}\int_\R\int_\R e^{it(x-y)}\left(T_ye^{-yc}\right)dy\,dt
=
e^{xc}\delta_x\left(T_ye^{-yc}\right)=T_x.
\end{eqnarray*}
}}
}
\end{rmk}
Now we assume $m=1$ and set $\sab(1)=:\sab$.
\begin{lem}
For all $\xi\in\R$ the function $e^{-\xi x}$ is a multiplier in  
$\sab$ ($\alpha<1$).
\end{lem}
\Pf
Let $\varphi\in\sab$. Then the function $f(z):=e^{-\xi z}\varphi(z)$ is 
entire and one has $$|f(z)|\leq C\exp\left(-a|x|^{\frac{1}{\alpha}}
+b|y|^{\frac{1}{1-\beta}}\right)e^{|\xi||x|}$$ which is lower than 
$C'\exp\left(-a'|x|^{\frac{1}{\alpha}}
+b|y|^{\frac{1}{1-\beta}}\right)$ as soon as $\alpha<1$.
\EPf
In particular, one has $\sab\subset\s'_\R=\cap_\Gamma\s'_\Gamma$ as soon as 
$\alpha<1$.
\begin{prop}
The Laplace transformation yields a linear isomorphism
$$
\LT:\left(\sab\right)_x\to J^*\sba
$$
where $J:\C\to\C$ denotes the multiplication by $i=\sqrt{-1}$.
\end{prop}
\Pf
Let $\varphi\in\sab$. Then 
$(J^*\LT(\varphi))(x+iy)=\int_\R e^{ixt-yt}\varphi(t)\,dt$. Hence 
$(J^*\LT(\varphi))|_\R(x)=(\FT(\varphi))(x)\in\sba$ (\cite{G}). Thus 
$\LT(\varphi)\in J^*\sba$. Now let $f\in J^*\sba$. One has 
$|f(x+iy)|\leq C\exp\left(-a|y|^{\frac{1}{\alpha}}
+b|x|^{\frac{1}{1-\beta}}\right)$ which guarantees that 
$f\in\o_\R$. Therefore
$\LT^{-1}f(x)=e^{xc}\int_\R e^{ixt}f(c+it)\,dt\quad(c\in\R)$. Choosing 
$c=0$, one gets $\LT^{-1}f=\FT J^*f\in\sab$. One therefore has an 
isomorphism $\LT^{-1}:J^*\sba\to\sab$.
\EPf
\section{Twisting maps}\label{TM}
\begin{dfn} Let $q\in\R$ and $\theta\in[0,2\pi[$. We define the twisting 
map $\phi_{q,\theta}:\C\to\C$ by 
$$
\left\{
\begin{array}{ccc}
\phi_{q,\theta}(z) & = & \frac{e^{i\theta}}{q}\sinh(iqz) \mbox{ if } q\neq 
0\\
\phi_{0,\theta}(z) & = & z.
\end{array}
\right.
$$
\end{dfn}
\begin{lem}\label{TWIST}
The twisting map $\phi_{q,0}\quad(q\neq 0)$ establishes a biholomorphic 
diffeomorphism 
$$
\phi_{q,0}:S_q:=\{z\in\C|\,|\Re{(z)}|<\frac{\pi}{2q}\}\to
\C-\{\pm i[\frac{1}{q},\infty[\}.
$$
\end{lem}
\Pf
In coordinates $z=x+iy$, the twisting map is
$$\phi_{q,0}(x,y)=\left(\frac{1}{q}\sinh(qy)\cos(qx),\frac{1}{q}\cosh(qy)\sin(qx)\right).$$
In particular, the imaginary ax\-is $x=0$ is sent onto the real ax\-is $y=0$, 
while the image of the vertical line $x=\pm\frac{\pi}{2q}$ under $\phi_{q,0}$ 
is the half imaginary line $\pm i[\frac{1}{q},\infty[$. At last, for 
$c\in]0,\frac{\pi}{2}[$ the image of the vertical line $x=\pm\frac{c}{q}$ 
is the branch of hyperbola $\{-\left(\frac{qx}{\cos 
c}\right)^2+\left(\frac{qy}{\sin c}\right)^2=1\}\cap\{\pm y>0\}$. For 
$\epsilon\in]0,\frac{\pi}{2}[$, one therefore gets a holomorphic 
diffeomorphism between the strip 
$]-\frac{\epsilon}{q},\frac{\epsilon}{q}[+i\R$ and the region 
$\{-\left(\frac{qx}{\cos 
\epsilon}\right)^2+\left(\frac{qy}{\sin \epsilon}\right)^2<1\}$.
\EPf
As explained in Section~\ref{FORMAL}, we are interested in considering 
transformations of the type $\LT^{-1}\circ\phi_{q,\theta}^*\circ\LT$ or 
$\LT^{-1}\circ(\phi_{q,\theta}^{-1})^*\circ\LT$. Let $F$ be a function 
defined on some domain $\Omega$ of $\C$. In order to define 
$\LT^{-1}(\phi_{q,\theta}^{-1})^*F$, we want 
$(\phi_{q,\theta}^{-1})^*F\in\o_\Gamma$. In particular we want $(\phi_{q,\theta}^{-1})^*F$
to be defined on vertical lines. This imposes to $\theta$ to be an integral 
multiple of $\frac{\pi}{2}$.
\begin{prop}\label{ISOTWIST}
Let $\alpha,\beta\in]0,1[$ be such that $\alpha+\beta\geq1$.
\begin{enumerate}
\item[(i)] For all open interval $I$ of positive numbers $(I\subset\R_0^+)$, one has 
the injection $(\phi_{q,0}^{-1})^*:J^*\sba\to\o_I\cap\o_{-I}$.
\item[(ii)] For all open interval 
$I\subset]-\frac{\pi}{2q},\frac{\pi}{2q}[$, one has the injection
$(\phi_{q,\frac{\pi}{2}}^{-1})^*:J^*\sba\to\o_I$.
\end{enumerate}
\end{prop}
\Pf
Let $F\in J^*\sba$ and set for simplicity $\phi=\sinh(iz)$. Let 
$I\subset]0,\infty[$ and $K$ be a compact set in $I$. Then $\phi^{-1}$ is 
well defined on the strip $K+i\R$ and $\phi^{-1}(K+i\R)\subset]-\frac{\pi}{2},
\frac{\pi}{2}[+i\R$ (cf. Lemma~\ref{TWIST}). Now, consider the integral
$I_K:=\int_{K+i\R}|(\phi^{-1})^*F(z)|\,|dz\wedge\overline{dz}|$. If 
$I_K<\infty$ for all $K$, then $(\phi^{-1})^*F\in\o_I$. Changing the 
variables $z=\phi(w)$, one gets 
$I_K=\int_{\phi^{-1}(K+i\R)}|F(w)|\,|\mbox{Jac}_\phi(w)|\,|dw\wedge\overline{dw}|$ 
where $\mbox{Jac}_\phi$ is the Jacobian determinant of $\phi$. A computation 
yields $\mbox{Jac}_\phi(w)=\cos^2(x)+\sinh^2(y)\quad(w=x+iy)$ and, since 
$\phi^{-1}(K+i\R)\subset]-\frac{\pi}{2},
\frac{\pi}{2}[+i\R$, one gets $I_K\leq\int^{\frac{\pi}{2}}_{-\frac{\pi}{2}}
\int^{\infty}_{-\infty}\exp\left(-a|y|^{\frac{1}{\beta}}
+b|x|^{\frac{1}{1-\alpha}}\right)\,\left(\cos^2(x)+\cosh^2(y)\right)\,dy\,dx$ 
which is lower than 
$2b\exp\left[\left(\frac{\pi}{2}\right)^{\frac{1}{1-\alpha}}\right]\int^{\infty}_{-\infty}
\exp\left(-a|y|^{\frac{1}{\beta}}\right)\,(1+\cosh^2(y))\,dy$, thus finite 
as soon as $\beta<1$. The exact same argument works for 
$I\subset]-\frac{\pi}{2},0[$, hence $(\phi^{-1})^*F\in\o_I\cap\o_{-I}$. A 
similar argument yields item (ii).
\EPf
\pagebreak
\section{Twisted Weyl product}
\subsection{The product formula}
Let us consider the fundamental space 
$\s^{\sigma(\alpha_1,\alpha_2)}_{(\alpha_1,\alpha_2)}(2)\qquad(\alpha_1,\alpha_2)\in\R^{2}$ 
(cf. Section~\ref{TYPE-S}). 
Let $\varphi\in\sA$ and consider the partial function 
$\varphi_{x_1}:x\mapsto\varphi(x_1,x)$. For all $x_1\in\R$, the function 
$\varphi_{x_1}$ belongs to ${\cal S}^{\alpha_{1}}_{\alpha_{2}}(1)=:{\cal S}^{\alpha_{1}}_{\alpha_{2}}$. Therefore provided some 
restrictions on $(\alpha_1,\alpha_2)$, the function 
$\LT^{-1}(\phi^{-1}_{q,k\frac{\pi}{2}})^*\LT(\varphi_{x_1})$ ($k=0,1$) is 
well defined as an element of $\s'_I$ (cf. Proposition~\ref{ISOTWIST}).
\begin{dfn}
We define the linear map 
$$
\sA\stackrel{\tau^{(k)}_q}{\longrightarrow}C^\infty(\R^2)\quad(k=0,1)
$$
by $$\tau^{(k)}_q:=\mbox{\rm id}_{x_1}\otimes\left(\LT^{-1}\circ
(\phi^{-1}_{q,k\frac{\pi}{2}})^*\circ\LT\right)_{x_2}\qquad(x_1,x_2)\in\R^2.$$
We denote by $E^{(k)}_{(\alpha_1,\alpha_2)}$ its range in $C^\infty(\R^2)$. The 
inverse map $\mbox{\rm id}_{x_1}\otimes\left(\LT^{-1}\circ
\phi_{q,k\frac{\pi}{2}}^*\circ\LT\right)_{x_2}|_{\Ekab}$ will be denoted 
by $\Tkq$. It yields a linear isomorphism $\Tkq:\Ekab\to\sA$.
\end{dfn}
\begin{prop}
The formula 
$$
u\skq v:=\taukq\left(\Tkq u\sw\Tkq v\right)
$$
defines an associative $\R$-algebra structure $\skq$ on $\Ekab$.
\end{prop}
\Pf
One has $\Tkq u,\Tkq v\in\sA$ provided $u,v\in\Ekab$. We know that $\sA$ is 
stable under Weyl's product $\sw$ (cf. Proposition~\ref{WEYL}). Hence 
$u\skq v$ is well defined as an element of $\Ekab$. The associativity is 
obvious since $\skq$ is nothing else than the transportation of Weyl's
product via the isomorphism $\Tkq:\Ekab\to\sA$.
\EPf
\begin{dfn}
The product $\skq$ on $\Ekab$ will be referred as the twisted Weyl product.
\end{dfn}
\subsection{Observables of exponential type}
Given a convergent star product, an important question is the one of 
existence of non-tempered observables in the domain of the star product. 
This question has been studied for the case of the Moyal-Weyl product in 
\cite{OMMY}.

Let $A=(\alpha_1,\alpha_2)\in]0,1[^2$ with $\alpha_{1}+\alpha_{2}\geq1$. Set $\s_A:=\sA$. 
Viewing $\s_A=\left(\s_A\right)_x\quad(x\in\R^2)$ as a subspace of 
$C^\infty(\R^2)$, we consider the following alternate presentation of 
$\s_A$. Consider the sequence of maps 
$$
\begin{array}{ccccc}
\left(\s_A\right)_x& \to&\o(\C^2)&\to&C^\infty(\R^2)\\
f(x_1,x_2)&\mapsto&f(z_1=x_1+iy_1, z_2=x_2+iy_2)&\mapsto&f(iy_1,x_2).
\end{array}
$$
The function $\hat{f}(y_1,x_2):=f(iy_1,x_2)$ determines completely $f$. So 
that we have an injection $\left(\s_A\right)_x\to 
C^\infty(\R^2):f\mapsto\hat{f}$. Remark that the space $\hat{\s_A}$ 
contains elements of exponential growth. For example, one has
$f=e^{-(z_1^2+z_2^2)}\in\s_{(\frac{1}{2},\frac{1}{2})}$, which yields 
$|\hat{f}(y_1,x_2)|=e^{y_1^2-x_2^2}$.

We denote by $\widehat{\sw}$ the product on $\hat{\s_A}$ obtained by 
transporting Weyl's product on $\left(\s_A\right)_x$ via $f\mapsto\hat{f}$.
Observe that $\hat{\s_A}$ is still stable under the pointwise 
multiplication whose $\widehat{\sw}$ is a non-commutative deformation of. 
Observe also that for every $\psi\in\hat{\s_A}$ and $y_1\in\R$ the 
partial function $x_2\mapsto\psi(y_1,x_2)$ is in ${\cal S}^{\alpha_{1}}_{\alpha_{2}}$. Therefore the 
transformations $\taukq$ and $\Tkq$ are well defined on $\hat{\s_A}$.
\begin{prop}
Set $\widehat{\Ekab}:=\taukq(\hat{\s_A})$. Then for all $a,b\in\widehat{\Ekab}$, 
the formula 
$$
a\;\widehat{\skq}\;b:=\taukq\left(\Tkq a\;\widehat{\sw}\;\Tkq b\right)
$$
defines an associative $\R$-algebra structure on $\widehat{\Ekab}$. The space 
$\widehat{\Ekab}$ contains elements of exponential growth.
\end{prop}

\end{document}